OpenStudy (anonymous):

Find the value of y when x equals -4 in this equation. 5x + (-3y) = 13

OpenStudy (compassionate):

-20 + (3y) = 13 Does that help?

OpenStudy (anonymous):

same as last time

OpenStudy (compassionate):

20 +(-20) + (3y) = 13 + 20 3y = 33 ... ?

OpenStudy (anonymous):

5^-4 + (-3y) = 13

OpenStudy (anonymous):

first do 5^-4

OpenStudy (compassionate):

Solymr, that would be 5 * 5 * 5 * 5 = 0.0016 We want 5 * -4

OpenStudy (anonymous):

well what im saying is 5 times -4 giving you -20

OpenStudy (anonymous):

-20 + (-3y) = 13

OpenStudy (anonymous):

then add 20 to both sides of the equations

OpenStudy (compassionate):

-3y 30 --- = --- = ? -3 -13

OpenStudy (anonymous):

no you divide -3 to both sides

OpenStudy (compassionate):

It's an obvious typo.

OpenStudy (anonymous):

-3y = 33 -------- -3 -3

OpenStudy (anonymous):

oh my B :PP lol

OpenStudy (anonymous):

okay what do i do?

OpenStudy (compassionate):

Divide: 33 --- -3

OpenStudy (anonymous):

so you get -11

OpenStudy (compassionate):

x = -?

OpenStudy (anonymous):

y = -11

OpenStudy (anonymous):

11

OpenStudy (anonymous):

hope it helped

OpenStudy (compassionate):

Good luck.

OpenStudy (anonymous):

is -11 fasure fasure right?

OpenStudy (compassionate):

-11 is correct.